OpenStudy (anonymous):

She earns 15$ a week plus 3$for each subscription she sells. How many subscriptions must she sell to earn 90$ a week

OpenStudy (anonymous):

Answer choices.. 30 25 40 or 100

OpenStudy (anonymous):

15*(3)x=90 Solve for X X=25

OpenStudy (anonymous):

25 * 3 = 75 75 + 15 = 90

OpenStudy (anonymous):

do you get it?

OpenStudy (anonymous):

So the answer is 25

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

ur welcome